# The star HIP 41378 potentially misaligned with its cohort of long-period planets

### v1

@@
The star HIP 41378 potentially misaligned with its cohort of long-period planets
@@ 2 Observations and data reduction
HARPS-N The ingress of the transit of HIP 41378 f was observed with the high-resolution (R115000), optical (383-693 nm) HARPS-N spectrograph (Cosentino et al. 2012) at the 3.6-m Telescopio Nazionale Galileo (TNG) at the Roque de Los Muchachos Observatory on the island of La Palma, Spain, as part of the programme ID OPT22B_15. Those observations were performed near the ingress time (between UTC 02:05 and 06:41). More spectra were collected the night following the transit event to extend the baseline. In total, we obtained 102 spectra of HIP 41378 with an exposure time of 300 s, leading to a Signal-to-noise (SNR) per pixel up to 70 at 550 nm. The observations were performed with the sky in the second fiber to monitor the sky background that would be important to remove for atmospheric study. The instrument is stable enough to allow precise RV constraints over one night. The spectra were reduced using the online pipeline, version 3.7, available at the observatory. This pipeline is based on the cross-correlation technique (Baranne et al. 1996) to compute an averaged line-profile and determine the instantaneous RV. This pipeline is routinely used in HARPS-N observations and publications.
@@
CARMENES The CARMENES spectrograph (R94600; Quirrenbach et al. 2018) at the 3.5m telescope of the Calar Alto Observatory in Spain also monitored the transit ingress contemporaneously with HARPS-N. The observations were obtained as part of the programme ID OPT22B_15. A total of 18 spectra were collected over the transit night with an exposure time of 900s, leading to a SNR up to 150 in the spectral order 50. A Fabry-Perot lamp illuminated the second fiber to monitor any instrumental drift within the transit night. Because the near-infrared arm of CARMENES was not stable enough and polluted by tellurics, we only used the visible arm to derive the RV (550-950 nm). The basic reduction of the spectra was carried out at the observatory with the CARACAL pipeline (Zechmeister et al. 2014; Bauer et al. 2015) version 2.20. RVs were subsequently extracted with the SHAQ pipeline, specifically developed for the KOBE experiment (Lillo-Box et al. 2022). This cross-correlation-based (Baranne et al. 1996) pipeline employs publicly available binary masks from the ESPRESSO instrument pipeline (Pepe et al. 2021). SHAQ has demonstrated compatibility with results obtained through template-matching algorithms (Balsalobre-Ruza et al. 2025).
@@
MaHPS We observed the ingress of the transit of HIP 41378 f with the MaHPS spectrograph installed on the 2.1m telescope at the Wendelstein Observatory. MaHPS consists of the fiber-fed, high-resolution (), optical ( nm) echelle spectrograph FOCES (Pfeiffer et al. 1998) and a Menlo Systems Laser Frequency Comb (LFC) as a wavelength calibration source. The observations lasted from 23:30 UT until 05:45 UT and covered a baseline before the transit and the ingress. The exposure time was 900 s for each observation, and we used simultaneous exposures of the LFC to correct for drifts during the observation. In total, we collected 22 spectra. At the beginning of the observation, the target was below 30 ∘ elevation. We, therefore, excluded the first 5 spectra from the analysis. The data was reduced with the GAMSE pipeline (Wang et al. 2016), and RVs were extracted using MARMOT (Kellermann 2021) following the analysis in Thomas et al. (2025).
@@
ESPRESSO The transit was also observed with the high-precision (R140000), visible (380-788 nm) ESPRESSO spectrograph (Pepe et al. 2021), mounted on the 8.2-m ESO-VLT (VLT) at Paranal Observatory in Chile, as part of the programme ID 104.20UB.004. Baseline observations were not secured as the transit occurred during a technical night near the middle of a 2-week interferometric campaign using the four VLTs. A total of 31 spectra (between UTC 06:25 and 09:12) were secured with an exposure time of 300 s, leading to a SNR per pixel at the level of 50, at 450 nm. The RVs were extracted using the ESPRESSO data-reduction software in its version 3.2.5.
@@
NEID On UT 13 November 2022 we observed HIP 41378 using the high-resolution () NEID optical spectrograph (380-930 nm) on the 3.5-m WIYN telescope at Kitt Peak National Observatory in AZ, USA (Schwab et al. 2016). We collected 25 spectra of HIP 41378 for 4 hours covering the second quarter of the transit of planet f (between UTC 08:33 and 11:54). Exposure times were fixed to 600 seconds, and the data were reduced with version 1.2.0 of the NEID Data Reduction Pipeline. This version includes an update to the calibration files which were required due to the thermal cycling of the NEID spectrometer after the Contreras fire. RVs and their uncertainties were obtained with the Cross-Correlation Function (CCF) method, resulting in measurement uncertainties of 3-4 m. In addition, we obtained six spectra respectively on UT nights 12 and 14 November 2022 to establish the out-of-transit baseline.
@@
EXPRES We observed HIP 41378 during a transit of planet f on UT night 13 November 2022 using the EXPRES spectrograph (Jurgenson et al. 2016) mounted on the 4.3-m Lowell Discovery Telescope (Levine et al. 2012; Levine & DeGroff 2016), located near Flagstaff, AZ, USA. EXPRES is a stabilized high-resolution () optical (390-780 nm) spectrograph with a photon noise precision of about 33 cm (Blackman et al. 2020). We took 16 exposures over a duration of hours covering the second quarter of the 19-hour duration transit (between UTC 09:01 and 13:15). Exposure times ranged from 600-900 seconds, resulting in a SNR between 43-69 at 550 nm. A LFC and Th-Ar lamp was used for wavelength calibration throughout the observations. We computed CCFs from the extracted 1D spectra using a F9 line mask. RVs and their measurement uncertainties were derived from the centre of a Gaussian fit to the CCF, resulting in a RV precision of 2-4 m. We also obtained 3 exposures in sequence on UT night 12 November 2022 and 7 exposures on UT night 14 November 2022 to establish the out-of-transit baseline.
@@
HIRES We acquired RV observations of HIP 41378 on 12 November 2022, 13 November 2022, and 14 November 2022 using the High Resolution Echelle Spectrometer (HIRES, (Vogt et al. 1994)) on the Keck I telescope at W. M. Keck Observatory. HIRES is an echelle spectrometer with a resolving power of 67,000, depending on the slit used, and a wavelength range spanning from 374 to 970 nm. We extracted RVs from the spectroscopic data (reduced using the standard procedure described in Howard et al. (2010)), using the iodine cell method outlined in Butler et al. (1996). Observations were taken using a 14” x 0.861” slit with exposure times determined by an internal exposure meter. The typical exposure lasted 4-5 minutes, resulting in a spectrum that reached a signal-to-noise ratio per resolution element of 200 near 550nm. The observations on 13 November 2022 were acquired directly before, during, and after the ingress of the transit of HIP 41378 f and covered approximately half of the transit (between UTC 11:42 and 15:40). The observations on 12 November 2022 and 14 November 2022 were acquired to help establish the out-of-transit RV trend.
@@
PARAS The end of the transit egress was visible by the PARAS spectrograph. PARAS (now known as PARAS-1) is a fibre-fed, high-resolution (R 67000), echelle spectrograph that is pressure and temperature-stabilized, operating within the wavelength range of 380 to 690 nm to facilitate precise RV measurements (Chakraborty et al. 2014). It is mounted on the PRL 1.2-m telescope located at the Gurushikhar observatory, Mt. Abu, India. The spectrograph has shown 1-3 m s-1 long-term RV stability on bright RV standard stars (Chakraborty et al. 2014; Sharma & Chakraborty 2021; Chakraborty et al. 2018). A total of six spectra of HIP 41378 were acquired using PARAS. These spectra were collected on November 12 and 13, 2022, utilising the simultaneous wavelength calibration mode with a Uranium-Argon (UAr) hollow cathode lamp (HCL), as outlined in Sharma & Chakraborty (2021). The exposure durations were set at 1200 seconds and 1800 seconds for the different spectra, leading to an SNR per pixel ranging from 13 to 19 at the blaze wavelength of 550 nm. Data reduction was performed using an automated, custom-designed pipeline developed in IDL (Chakraborty et al. 2014), based on the algorithms described by Piskunov & Valenti (2002). The resulting extracted spectra from this reduction process were subsequently used for RV measurements. RVs were obtained by cross-correlating the stellar spectra with a numerical template mask representative of the same spectral type (Baranne et al. 1996; Pepe et al. 2002). The errors provided here are based on the fitting errors of the CCF and the photon noise, computed similarly to the methods presented in Chaturvedi et al. (2016, 2018).
@@
HERMES 23 spectra were obtained with the high-resolution (R85000) fibre-fed spectrograph HERMES mounted on the Flemish 1.2m Mercator telescope on the Roque de los Muchachos observatory in Spain (Raskin et al. 2011). This instrument is equipped with a science-grade pipeline which returns the response corrected, wavelength-calibrated 1D full spectrum from about 390 to 900nm on the barycentric velocity frame. 10 spectra were collected at ingress time (between UTC 02:52 and 06:40) and 12 spectra after the transit.
@@
K2 short-cadence photometry HIP 41378 was initially discovered and subsequently re-observed by the Kepler space telescope during the K2 extended mission, with data collected in Campaign 5 (2015; Vanderburg et al. 2016) and Campaign 18 (2018; Berardo et al. 2019; Becker et al. 2019). On both occasions, a single transit of HIP 41378 f was recorded. The Campaign 18 observations were obtained in short-cadence mode, offering improved temporal resolution. We incorporated this high-cadence K2 photometry into our analysis, alongside the RM effect detection, to refine the planetary parameters and constrain the system’s true spin-orbit obliquity. The short-cadence data from Campaign 18 were processed using the K2SFF pipeline (Vanderburg & Johnson 2014), which corrects for the telescope’s pointing drifts. To preserve the transit shape and depth while mitigating systematics, we applied a self-flat-fielding approach that explicitly accounts for the presence of the transit. Remaining instrumental or stellar variability was modelled and removed using a cubic spline with knots spaced every 10 days, resulting in a cleaner, transit-preserving light curve suitable for joint modelling with the RM observations.
@@ 3.1 Stellar rotation
The rotational modulation of HIP 41378 is clearly visible in the K2 light curve. We used the short-cadence Campaign 18 light curve, reduced with the EVEREST pipeline, to estimate the stellar rotation period. The transits of planets b, c, d, and f were masked using the ephemerides and transit durations from Becker et al. (2019). An initial estimate with a Lomb-Scargle periodogram (Zechmeister & Kürster 2009) gives a rotation period of days (see figure 7). To confirm this result, we also applied a Bayesian Generalised Lomb-Scargle periodogram (Mortier et al. 2015), which revealed a single prominent peak at approximately 6.4 days. Finally, an autocorrelation function (ACF) analysis following McQuillan et al. (2013) indicated a similar period of roughly 6.7 days (see Figure 7), consistent with the previous estimates. To further refine the stellar rotation period, we applied quasi-periodic Gaussian Process (GP) modeling using the celerite2 package, which implements a kernel composed of two simple harmonic oscillators centered at and (Angus et al. 2018). A uniform prior was adopted for the rotation period over the range 4–20 days. The GP model yields an inferred rotation period of days.
In addition to photometry, the radial-velocity (RV) analysis of ESPRESSO data (Grouffal et al. 2025 in prep.) revealed a stellar rotation period of days, as inferred from the index, along with a signal in the periodogram near 8 days. To be as conservative as possible, all available data suggest that the stellar rotation period lies between 6 and 9 days. We therefore adopted a uniform prior between 6 and 9 days for the rotation period of the star in the following analysis.
@@ 3.2 Planetary system parameters
In starry, the adopted geometry of the system is such that the plane is the plane of the sky and the axis points toward the observer. The planet orbits around the star in a counter-clockwise manner. In this context, the projected obliquity is defined as the angle between the stellar rotation axis and the angular momentum vector of the planet projected on the sky plane (Montet et al. 2020). This convention differs from other RM models like ARoME (Boué et al. 2013) and can induce opposite solutions for the value of .
@@ 3.4 Macroturbulence
With ARoME, we obtain a projected obliquity of ∘ and a projected rotational velocity of km s-1. The three-dimensional obliquity is ∘. The fit converges toward a lower macroturbulence value, km s-1, than the one predicted by the relation from (Doyle et al. 2014). Forcing the macroturbulence to that empirical value results in a poor fit to the data. The best fit is represented in Figure 10, and the parameters and priors are in Table 10.
@@ 3.5 Rossiter-McLaughlin Revolutions
The HARPS-N data allow for two possible configurations: one with the stellar north pole visible (∘) and positive , and the other with the south pole visible (∘) and positive . In both cases, the best-fit favours high values for , reinforcing the findings from the classical analysis. While the HARPS-N data alone favours the southern configuration, we isolate the northern one for consistency with the classical analysis of all datasets. It corresponds to ∘, ∘, and at the 1 level. Combining the PDF for , , and the orbital inclination yields a 3D spin-orbit angle ∘.

### v2

# Key Words.: Planets and satellites: individual: HIP41378, Techniques: spectroscopic, Stars: rotation1 Introduction

@@
The star HIP 41378 potentially misaligned with its cohort of long-period planets ††thanks: The radial velocity data for all instruments are available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ 2 Observations and data reduction
The ingress of the transit of HIP 41378 f was observed with the high-resolution (R115000), optical (383-693 nm) HARPS-N spectrograph (Cosentino et al. 2012) at the 3.6-m Telescopio Nazionale Galileo (TNG) at the Roque de Los Muchachos Observatory on the island of La Palma, Spain, as part of the programme ID OPT22B_15. Those observations were performed near the ingress time (between UTC 02:05 and 06:41). More spectra were collected the night following the transit event to extend the baseline. In total, we obtained 102 spectra of HIP 41378 with an exposure time of 300 s, leading to a Signal-to-noise (SNR) per pixel up to 70 at 550 nm. The observations were performed with the sky in the second fibre to monitor the sky background that would be important to remove for atmospheric study. The instrument is stable enough to allow precise RV constraints over one night. The spectra were reduced using the online pipeline, version 3.7, available at the observatory. This pipeline is based on the cross-correlation technique (Baranne et al. 1996) to compute an averaged line-profile and determine the instantaneous RV. This pipeline is routinely used in HARPS-N observations and publications.
CARMENES
@@
The CARMENES spectrograph (R94600; Quirrenbach et al. 2018) at the 3.5m telescope of the Calar Alto Observatory in Spain also monitored the transit ingress contemporaneously with HARPS-N. The observations were obtained as part of the programme ID OPT22B_15. A total of 18 spectra were collected over the transit night with an exposure time of 900s, leading to a SNR up to 150 in the spectral order 50. A Fabry-Perot lamp illuminated the second fibre to monitor any instrumental drift within the transit night. Because the near-infrared arm of CARMENES was not stable enough and polluted by tellurics, we only used the visible arm to derive the RV (550-950 nm). The basic reduction of the spectra was carried out at the observatory with the CARACAL pipeline (Zechmeister et al. 2014; Bauer et al. 2015) version 2.20. RVs were subsequently extracted with the SHAQ pipeline, specifically developed for the KOBE experiment (Lillo-Box et al. 2022). This cross-correlation-based (Baranne et al. 1996) pipeline employs publicly available binary masks from the ESPRESSO instrument pipeline (Pepe et al. 2021). SHAQ has demonstrated compatibility with results obtained through template-matching algorithms (Balsalobre-Ruza et al. 2025).
MaHPS
@@
We observed the ingress of the transit of HIP 41378 f with the MaHPS spectrograph installed on the 2.1m telescope at the Wendelstein Observatory. MaHPS consists of there-fed, high-resolution (), optical ( nm) echelle spectrograph FOCES (Pfeiffer et al. 1998) and a Menlo Systems Laser Frequency Comb (LFC) as a wavelength calibration source. The observations lasted from 23:30 UT until 05:45 UT and covered a baseline before the transit and the ingress. The exposure time was 900 s for each observation, and we used simultaneous exposures of the LFC to correct for drifts during the observation. In total, we collected 22 spectra. At the beginning of the observation, the target was below 30 ∘ elevation. We, therefore, excluded the first five spectra from the analysis. The data was reduced with the GAMSE pipeline (Wang et al. 2016), and RVs were extracted using MARMOT (Kellermann 2021) following the analysis in Thomas et al. (2025).
ESPRESSO
@@
The transit was also observed with the high-precision (R140000), visible (380-788 nm) ESPRESSO spectrograph (Pepe et al. 2021), mounted on the 8.2-m ESO-VLT (VLT) at Paranal Observatory in Chile, as part of the programme ID 104.20UB.004. Baseline observations were not secured as the transit occurred during a technical night near the middle of a 2-week interferometric campaign using the four VLTs. A total of 31 spectra (between UTC 06:25 and 09:12) were secured with an exposure time of 300 s, leading to a SNR per pixel at the level of 50, at 450 nm. The RVs were extracted using the ESPRESSO data-reduction software in its version 3.2.5.
NEID
@@
On UT 13 November 2022, we observed HIP 41378 using the high-resolution () NEID optical spectrograph (380-930 nm) on the 3.5-m WIYN telescope at Kitt Peak National Observatory in AZ, USA (Schwab et al. 2016). We collected 25 spectra of HIP 41378 for 4 hours covering the second quarter of the transit of planet f (between UTC 08:33 and 11:54). Exposure times were fixed to 600 seconds, and the data were reduced with version 1.2.0 of the NEID Data Reduction Pipeline. This version includes an update to the calibration files, which were required due to the thermal cycling of the NEID spectrometer after the Contreras fire. RVs and their uncertainties were obtained with the Cross-Correlation Function (CCF) method, resulting in measurement uncertainties of 3-4 m. In addition, we obtained six spectra respectively on UT nights 12 and 14 November 2022 to establish the out-of-transit baseline.
EXPRES
@@
We observed HIP 41378 during a transit of planet f on UT night 13 November 2022 using the EXPRES spectrograph (Jurgenson et al. 2016) mounted on the 4.3-m Lowell Discovery Telescope (Levine et al. 2012; Levine & DeGroff 2016), located near Flagstaff, AZ, USA. EXPRES is a stabilised high-resolution () optical (390-780 nm) spectrograph with a photon noise precision of about 33 cm. We took 16 exposures over a duration of hours covering the second quarter of the 19-hour duration transit (between UTC 09:01 and 13:15). Exposure times ranged from 600-900 seconds, resulting in a SNR between 43-69 at 550 nm. A LFC and Th-Ar lamp was used for wavelength calibration throughout the observations. We computed CCFs from the extracted 1D spectra using a F9 line mask. RVs and their measurement uncertainties were derived from the centre of a Gaussian fit to the CCF, resulting in a RV precision of 2-4 m. We also obtained three exposures in sequence on UT night 12 November 2022 and 7 exposures on UT night 14 November 2022 to establish the out-of-transit baseline.
HIRES
@@
We acquired RV observations of HIP 41378 on 12 November 2022, 13 November 2022, and 14 November 2022 using the High Resolution Echelle Spectrometer (HIRES, (Vogt et al. 1994)) on the Keck I telescope at W. M. Keck Observatory. HIRES is an echelle spectrometer with a resolving power of 67,000, depending on the slit used, and a wavelength range spanning from 374 to 970 nm. We extracted RVs from the spectroscopic data (reduced using the standard procedure described in Howard et al. (2010)), using the iodine cell method outlined in Butler et al. (1996). Observations were taken using a 14” x 0.861” slit with exposure times determined by an internal exposure meter. The typical exposure lasted 4-5 minutes, resulting in a spectrum that reached a signal-to-noise ratio per resolution element of 200 near 550nm. The observations on 13 November 2022 were acquired directly before, during, and after the ingress of the transit of HIP 41378 f and covered approximately half of the transit (between UTC 11:42 and 15:40). The observations on 12 November 2022 and 14 November 2022 were acquired to help establish the out-of-transit RV trend.
PARAS
@@
The end of the transit egress was visible by the PARAS spectrograph. PARAS (now known as PARAS-1) is a fibre-fed, high-resolution (R 67000), echelle spectrograph that is pressure and temperature-stabilised, operating within the wavelength range of 380 to 690 nm to facilitate precise RV measurements (Chakraborty et al. 2014). It is mounted on the PRL 1.2-m telescope located at the Gurushikhar observatory, Mt. Abu, India. The spectrograph has shown 1-3 m s-1 long-term RV stability on bright RV standard stars (Chakraborty et al. 2014; Sharma & Chakraborty 2021; Chakraborty et al. 2018). A total of six spectra of HIP 41378 were acquired using PARAS. These spectra were collected on November 12 and 13, 2022, utilising the simultaneous wavelength calibration mode with a Uranium-Argon (UAr) hollow cathode lamp (HCL), as outlined in Sharma & Chakraborty (2021). The exposure durations were set at 1200 seconds and 1800 seconds for the different spectra, leading to an SNR per pixel ranging from 13 to 19 at the blaze wavelength of 550 nm. Data reduction was performed using an automated, custom-designed pipeline developed in IDL (Chakraborty et al. 2014), based on the algorithms described by Piskunov & Valenti (2002). The resulting extracted spectra from this reduction process were subsequently used for RV measurements. RVs were obtained by cross-correlating the stellar spectra with a numerical template mask representative of the same spectral type (Baranne et al. 1996; Pepe et al. 2002). The errors provided here are based on the fitting errors of the CCF and the photon noise, computed similarly to the methods presented in Chaturvedi et al. (2016, 2018).
HERMES
@@
23 spectra were obtained with the high-resolution (R85000) fibre-fed spectrograph HERMES mounted on the Flemish 1.2m Mercator telescope on the Roque de los Muchachos observatory in Spain (Raskin et al. 2011). This instrument is equipped with a science-grade pipeline which returns the response corrected, wavelength-calibrated 1D full spectrum from about 390 to 900nm on the barycentric velocity frame. Ten spectra were collected at ingress time (between UTC 02:52 and 06:40) and 12 spectra after the transit.
K2 short-cadence photometry
@@
HIP 41378 was initially discovered and subsequently re-observed by the Kepler space telescope during the K2 extended mission, with data collected in Campaign 5 (2015; Vanderburg et al. 2016) and Campaign 18 (2018; Berardo et al. 2019; Becker et al. 2019). On both occasions, a single transit of HIP 41378 f was recorded. The Campaign 18 observations were obtained in short-cadence mode, offering improved temporal resolution. We incorporated this high-cadence K2 photometry into our analysis, alongside the RM effect detection, to refine the planetary parameters and constrain the system’s true spin-orbit obliquity. The short-cadence data from Campaign 18 were processed using the K2SFF pipeline (Vanderburg & Johnson 2014), which corrects for the telescope’s pointing drifts. To preserve the transit shape and depth while mitigating systematics, we applied a self-flat-fielding approach that explicitly accounts for the presence of the transit. Remaining instrumental or stellar variability was modelled and removed using a cubic spline with knots spaced every 10 days, resulting in a cleaner, transit-preserving light curve suitable for joint modelling with the RM observations.
@@ 3.1 Stellar rotation
The rotational modulation of HIP 41378 is clearly visible in the K2 light curve. We used the short-cadence Campaign 18 light curve, reduced with the EVEREST pipeline, to estimate the stellar rotation period. The transits of planets b, c, d, and f were masked using the ephemerides and transit durations from Becker et al. (2019). An initial estimate with a Lomb-Scargle periodogram (Zechmeister & Kürster 2009) gives a rotation period of days (see figure 7). To confirm this result, we also applied a Bayesian Generalised Lomb-Scargle periodogram (Mortier et al. 2015), which revealed a single prominent peak at approximately 6.4 days. Finally, an autocorrelation function (ACF) analysis following McQuillan et al. (2013) indicated a similar period of roughly 6.7 days (see Figure 7), consistent with the previous estimates. To further refine the stellar rotation period, we applied quasi-periodic Gaussian Process (GP) modelling using the celerite2 package, which implements a kernel composed of two simple harmonic oscillators centred at and (Angus et al. 2018). A uniform prior was adopted for the rotation period over the range 4–20 days. The GP model yields an inferred rotation period of days.
In addition to photometry, the radial-velocity (RV) analysis of ESPRESSO data (Grouffal et al. in prep.) revealed a stellar rotation period of days, as inferred from the index, along with a signal in the periodogram near 8 days. To be as conservative as possible, all available data suggest that the stellar rotation period lies between 6 and 9 days. We therefore adopted a uniform prior between 6 and 9 days for the rotation period of the star in the following analysis.
@@ 3.2 Planetary system parameters
In starry, the adopted geometry of the system is such that the plane is the plane of the sky and the axis points toward the observer. The planet orbits around the star in a counterclockwise manner. In this context, the projected obliquity is defined as the angle between the stellar rotation axis and the angular momentum vector of the planet projected on the sky plane (Montet et al. 2020). This convention differs from other RM models like ARoME (Boué et al. 2013) and can induce opposite solutions for the value of .
@@ 3.4 Macroturbulence
With ARoME, we obtain a projected obliquity of ∘ and a projected rotational velocity of km s-1. The three-dimensional obliquity is ∘. The fit converges toward a lower macroturbulence value, km s-1, than the one predicted by the relation from Doyle et al. (2014). Forcing the macroturbulence to that empirical value results in a poor fit to the data. The best fit is represented in Figure 10, and the parameters and priors are in Table 10.
@@ 3.5 Rossiter-McLaughlin Revolutions
The HARPS-N data allow for two possible configurations: one with the stellar north pole visible (∘) and positive , and the other with the south pole visible (∘) and positive . In both cases, the best-fit favours high values for , reinforcing the findings from the classical analysis. While the HARPS-N data alone favour the southern configuration, we isolate the northern one for consistency with the classical analysis of all datasets. It corresponds to ∘, ∘, and at the 1 level. Combining the PDF for , , and the orbital inclination yields a 3D spin-orbit angle ∘.